OpenStudy (anonymous):

SIMPLIFY: x^2 + x - 6 over 16 - 4x^2

OpenStudy (anonymous):

(x +3)(x-2) / (4-2x)(4+2x) =(x +3)(x-2) / 4(2-x)(2+x) =(x +3)(x-2) / -4(x-2)(2+x) (x-2) cancels =- (x+3) / 4(2+x)

OpenStudy (anonymous):

THANK YOU SO MUCH FOR THE EXPLANATION ! ;-)

OpenStudy (anonymous):

no probs